OpenStudy (johnnydicamillo):

Use implicit differentiation to find the equation of the tangent line to the curve xy^3+xy=10 at the point (5,1). The equation of this tangent line can be written in the form y=mx+b where m is:

OpenStudy (johnnydicamillo):

I need to find m and B

OpenStudy (johnnydicamillo):

So lets begin \[3y^2*y' + y' = 0\] I think that's how you simplify it.

OpenStudy (anonymous):

Nope. Gonna have to restart.

OpenStudy (johnnydicamillo):

wait, product rule?

OpenStudy (anonymous):

So, let's start with the first term just so we can get the hang of what we need to do. \(xy^{3}\). We have two variable expressions being multiplied, so we need to do product rule actually. Recall the product rule for derivatives is \(f'(x)g(x) + f(x)g'(x)\). So if we want to break this first one down, we can say: \(f(x) = x\), \(f'(x) = 1\) \(g(x) = y^{3}\), \(g'(x) = 3y^{2}*y'\) Now, I only tack on the \(y'\) when I actually do a derivative operation to "y". So using product rule, the first term becomes: \(y^{3} + x(3y^{2})y'\) Does the derivative for that first term make sense?

OpenStudy (johnnydicamillo):

okay, I believe so

OpenStudy (anonymous):

Alrighty, cool. So how do you think the derivative of the 2nd term would turn out then? It's product rule again :)

OpenStudy (johnnydicamillo):

\[(x) (y)*y' + (y)(1) * y'\]

OpenStudy (johnnydicamillo):

scratch the last y

OpenStudy (anonymous):

You only have a \(y'\) when you take the derivative of a y. Just because a term has a y in it doesn't mean you will have \(y'\) in it after you take the derivative. So if I do the derivative of xy. \(f(x) = x\), \(f'(x) = 1\) \(g(x) = y\), \(g'(x) = 1*y'\) Only that one derivative will have a \(y'\) in it. So that means I would have: \(y + x*y'\) Make sure that makes sense to ya.

OpenStudy (johnnydicamillo):

okay, now what do we do next

OpenStudy (anonymous):

Do you see how I did that, though?

OpenStudy (johnnydicamillo):

yes, implicit differentiation and product rule

OpenStudy (johnnydicamillo):

okay, now what about those extra ys

OpenStudy (johnnydicamillo):

okay so, it looks like we did something wrong here

OpenStudy (johnnydicamillo):

When I try to enter the slop online, it says its incorrect.

OpenStudy (anonymous):

I wrote it out incorrect as I was going through the steps. I missed a term, so that's on me :)

OpenStudy (anonymous):

Okay, so I should have: \(y^{3} + (3xy^{2})y' + y + (x)y'=0\) \((3xy^{2})y' + (x)y' = -y-y^{3}\) \(y'(3xy^{2} + x) = -y-y^{3}\) \(y' = \frac{-y-y^{3}}{3xy^{2} + x}\)

OpenStudy (johnnydicamillo):

okay, so we plug in "1" for y correct

OpenStudy (anonymous):

and 5 for x :3

OpenStudy (johnnydicamillo):

y' = -1/10

OpenStudy (johnnydicamillo):

and y-1 = -(1/10)x + 1/2 correct?

OpenStudy (anonymous):

yes

OpenStudy (johnnydicamillo):

so how to make this in y = mx + b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(y - 1 = \frac{-1}{10}(x-5)\) \(y = \frac{-x}{10} + \frac{1}{2} + 1\) \(y = \frac{-x}{10} + \frac{3}{2}\) That should be your answer :P TO get into y = mx + b, you just solve for y using your point-slope form equation.

OpenStudy (johnnydicamillo):

WE DID IT!!!!!

OpenStudy (johnnydicamillo):

Thanks!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Thanks! as thx